# SuperiorGAT: Graph Attention Networks for Sparse LiDAR Point Cloud Reconstruction in Autonomous Systems

### v2

@@ Abstract
LiDAR-based perception in autonomous systems is constrained by fixed vertical beam resolution and further compromised by beam dropout resulting from environmental occlusions. This paper introduces SuperiorGAT, a graph attention-based framework designed to reconstruct missing elevation information in sparse LiDAR point clouds. By modeling LiDAR scans as beam-aware graphs and incorporating gated residual fusion with feed-forward refinement, SuperiorGAT enables accurate reconstruction without increasing network depth. To evaluate performance, structured beam dropout is simulated by removing every fourth vertical scanning beam. Extensive experiments across diverse KITTI environments, including Person, Road, Campus, and City sequences, demonstrate that SuperiorGAT consistently achieves lower reconstruction error and improved geometric consistency compared to PointNet-based models and deeper GAT baselines. Qualitative X–Z projections further confirm the model’s ability to preserve structural integrity with minimal vertical distortion. These results suggest that architectural refinement offers a computationally efficient method for improving LiDAR resolution without requiring additional sensor hardware providing a foundation for real-time optimization on embedded hardware.
@@ Introduction
Autonomous vehicles rely critically on Light Detection and Ranging (LiDAR) sensors for 3D environmental perception, enabling precise object detection, localization, and mapping essential for safe navigation [zhang2022, geiger2013]. Despite their fundamental importance, LiDAR systems face a critical vulnerability: performance severely degrades under structured data loss, particularly beam dropout caused by hardware malfunctions or environmental occlusions, which disrupts spatial continuity and compromises downstream tasks such as object detection and path planning [zhang2022, behley2019]. This challenge is particularly acute for cost-effective LiDAR systems that employ fewer scanning beams (typically 16-32 beams compared to 64+ in premium systems), inherently producing sparse point clouds that demand robust reconstruction methods to achieve the high-resolution performance required for autonomous operation [dosovitskiy2017].
@@
Current reconstruction approaches face a fundamental trade-off between accuracy and computational efficiency, limiting their practical deployment. Compressive sensing (CS) methods provide theoretical guarantees for sparse signal recovery but depend on computationally intensive iterative solvers that render them impractical for the real-time constraints of autonomous systems [candes2006, donoho2006]. Convolutional neural networks (CNNs), particularly voxel-based architectures, demonstrate effectiveness in processing point clouds but necessitate dense 3D projections, resulting in inefficient memory usage and computational overhead when applied to inherently sparse LiDAR data [wang2018, hatun2023].
@@
Recent advances in deep learning, including hierarchical point cloud models that operate directly on unordered point sets, have shown promise but remain inadequate for addressing the structured sparsity patterns characteristic of beam dropout scenarios [qi2017]. While Graph Neural Networks (GNNs), especially Graph Attention Networks (GATs), have demonstrated significant potential in modeling complex geometric relationships within irregular data structures through sophisticated graph representations [scarselli2009, velickovic2018, turati2022], standard graph-based approaches, including graph convolutional networks, often fail to effectively prioritize the critical local geometric features essential for high-fidelity reconstruction of sparse LiDAR data [kipf2017]. Our recent conference work demonstrated that multi-layer Graph Attention Networks can effectively reconstruct missing LiDAR elevation values using raw point-cloud geometry alone, without relying on RGB images or temporal information [awedat2025]. While increasing network depth improved reconstruction fidelity, deeper attention stacks introduced higher computational cost and stability limitations. In this work, which extends our previous findings, we move beyond depth-based improvements and focus on architectural refinement to achieve superior performance with a lightweight, single-layer design incorporating gated residual fusion..
@@
To address these fundamental limitations, we propose SuperiorGAT, a novel graph attention-based framework specifically engineered to reconstruct sparse LiDAR point clouds with exceptional fidelity while maintaining the computational efficiency necessary for real-time deployment in autonomous systems [hatun2023, turati2022]. Our approach directly tackles the structured sparsity challenge through several key innovations. SuperiorGAT implements a realistic beam dropout simulation framework that accurately mimics hardware faults observed in real-world LiDAR deployments, providing a robust testing environment for autonomous vehicle applications [zhang2022, behley2019]. The system constructs optimized k-nearest neighbor graphs to effectively represent sparse point cloud structures while incorporating specialized beam index features that enhance both spatial and structural contextual understanding [turati2022].
@@
The model architecture features a carefully designed, lightweight framework that incorporates multiple graph attention layers with multi-head attention mechanisms and residual connections, enabling highly efficient processing of sparse data without the computational burden of iterative CS optimization or the memory overhead of dense CNN projections [velickovic2018]. By directly addressing structured sparsity patterns, SuperiorGAT provides a scalable and practical solution for autonomous driving applications, successfully bridging the critical gap between high-resolution reconstruction accuracy and real-time computational efficiency. Our primary contributions to the field include:
@@
Light Detection and Ranging (LiDAR) technology is a cornerstone of 3D perception for autonomous vehicles, enabling tasks like object detection and mapping through precise geometric data [geiger2013, behley2019]. Yet, sparse point clouds—caused by low-resolution sensors or beam dropout from environmental factors such as rain or hardware faults—complicate accurate reconstruction [zhang2021]. As illustrated in Figure 1, beam dropout predominantly disrupts elevation (z-coordinates), while x and y coordinates, derived from azimuthal scans, remain stable, underscoring the need for targeted z-reconstruction methods [eskandar2023, glennie2013]. This vertical sparsity arises from the limited number of laser channels in affordable LiDAR systems, a challenge evident in datasets such as KITTI [geiger2013].
@@
Conventional reconstruction techniques include interpolation and compressive sensing (CS). Interpolation methods, such as linear or nearest-neighbor approaches, estimate missing points based on proximity but often smooth out critical geometric features, particularly in complex scenes [candes2006, donoho2006, rusu2009]. Advanced techniques, such as spline interpolation, reduce this issue but can introduce artifacts in regions with sharp elevation shifts [rusu2009]. CS offers a theoretical framework for recovering sparse signals but demands computationally intensive iterative processes, rendering it unsuitable for real-time applications where sub-100 millisecond (ms) inference is critical [candes2006, donoho2006, velickovic2018]. These limitations underscore the need for efficient solutions specifically designed for LiDAR’s structured sparsity.
@@
Deep learning has transformed point cloud processing, with Convolutional Neural Networks (CNNs) like VoxelNet converting sparse data into dense voxel grids for tasks such as 3D object detection [wang2018]. However, this densification escalates computational costs and erodes fine details in sparse LiDAR data [wang2018, you2023]. VoxelNet’s cubic scaling with resolution contrasts with PointNet++, which employs hierarchical sampling to handle irregular data; yet, both methods falter with structured dropout patterns due to static feature aggregation [qi2017, graham2018]. PointNet and its extensions leverage permutation invariance but struggle to adapt to the specific sparsity of beam dropout [qi2017, eskandar2023].
@@
Recent advances in LiDAR super-resolution target vertical resolution enhancement, focusing on z-coordinates for improved pedestrian detection [you2023, eskandar2023]. You and Kim (2023) introduced a 2D range image-based upsampling technique, interpolating x, y, z, and intensity to boost 3D detection accuracy, while Eskandar et al. (2023) proposed HALS, a height-aware model using polar coordinate regression to excel on KITTI data [you2023, eskandar2023]. These approaches affirm z-reconstruction’s importance, given the reliability of x and y coordinates in LiDAR scans [zhang2021], though their reliance on dense architectures limits real-time viability.
Graph Neural Networks (GNNs) provide a versatile framework for modeling irregular data, where points are represented as nodes and relationships are represented as edges [scarselli2009]. Graph Convolutional Networks (GCNs) employ fixed-weight convolutions, which are effective for node classification but less adept at prioritizing sparse LiDAR geometries [kipf2017]. Enhanced GNNs incorporate edge attributes, such as distances, thereby enhancing spatial modeling [wang2019, turati2022]. Graph Attention Networks (GATs), pioneered by Veličković et al., introduce dynamic weighting via attention mechanisms, drawing from insights from neural machine translation [velickovic2018, bahdanau2015]. Although successful in network analysis, GATs remain underexplored for LiDAR, where their adaptive aggregation is particularly suited to sparse geometries [wang2019].
@@
Recent work in graph neural networks has explored architectural elements that enhance stability and expressive capacity beyond the standard GAT formulation. In particular, gated skip connections have been shown to improve gradient flow and help models balance raw node features with attention-refined representations [li2019deepgcns, bresson2017residual]. Building upon our prior work[awedat2025] , which established the effectiveness of multi-layer GAT architectures for recovering missing elevation data, this research extends that foundation by introducing a more efficient architectural design. By incorporating gated residual fusion and feed-forward refinement, we achieve improved reconstruction accuracy while eliminating the need for increasing graph depth.. Likewise, several graph transformer variants incorporate feed-forward sublayers similar to those used in sequence transformers to increase nonlinear capacity and improve local geometric modeling [dwivedi2020generalization]. These ideas motivate our adoption of a lightweight gated residual pathway and a transformer-style feed-forward block, adapted specifically to the characteristics of sparse LiDAR reconstruction.
@@
SuperiorGAT advances GATs by optimizing for LiDAR z-reconstruction, integrating beam index features to reflect scanning patterns, and employing a lightweight design with multi-head attention and residuals [velickovic2018, wang2019]. Unlike previous methods, it tackles structured beam dropout (Figure 1) with a focus on z-coordinates, offering a balance of accuracy and efficiency for autonomous driving. This sets it apart from dense CNN approaches [wang2018], PointNet variants [qi2017], and super-resolution techniques [you2023, eskandar2023], establishing a novel real-time solution.
@@
Graph Neural Networks (GNNs) represent a powerful class of deep learning models designed to perform inference on data described by graphs. By leveraging the inherent relational structure of graph data, GNNs overcome a key limitation of traditional neural networks, which require input data to exist in an independent and identically distributed (i.i.d.) or grid-like structure (e.g., images) [scarselli2009, wu2021survey].
The core operational principle of GNNs is message passing, a framework where nodes in a graph iteratively aggregate information from their local neighbors to build increasingly sophisticated representations of themselves and their context [gilmer2017neural, wu2021survey]. This process can be broken down into three steps at each layer $l$:
@@
This mechanism allows each node to progressively incorporate information from multi-hop neighborhoods. For example, in a 2-layer GNN, a central node aggregates features from its direct neighbors in the first layer, and from its neighbors’ neighbors in the second layer. This ability to capture higher-order structural dependencies makes GNNs particularly effective for tasks where node states are influenced by spatial and relational context [kipf2017, velickovic2018gat, wu2019simplifying].
@@
This graph structure enables the application of Graph Neural Networks (GNNs) to model local relationships among points, leveraging the beam-aware connectivity to infer missing z-values. The representation provides a foundation for advanced graph-based techniques, setting the stage for exploring convolutional and attention-based approaches to enhance reconstruction accuracy. This prepares the framework for the subsequent development of tailored GNN architectures.
@@
Graph Attention Networks (GATs) represent a significant advancement in graph representation learning by introducing an attention mechanism that enables dynamic, content-aware neighborhood aggregation [velickovic2018]. Unlike Graph Convolutional Networks (GCNs) that employ fixed-weight aggregation based on graph structure, GATs compute adaptive attention coefficients that weight the contribution of each neighbor during feature propagation, allowing the model to focus on the most relevant contextual information for sparse LiDAR reconstruction.
@@
$\alpha_{ij}=\frac{e^{(e_{ij})}}{\sum_{k\in\mathcal{N}(i)}e^{(e_{ik})}}$ (6)
@@
where $\sigma$ is a nonlinear activation function. This formulation enables anisotropic aggregation, where each neighbor’s contribution is dynamically weighted based on feature compatibility, providing a more expressive scheme than isotropic approaches [velickovic2018, gilmer2017neural]. The standard GAT architecture follows the sequential processing pipeline illustrated in Figure 3a.
@@
The core of our proposed framework is the SuperiorGAT layer, which is designed to capture complex local geometries without the computational burden of deep network stacks. While standard GATs as depicted in Figure 3(a) provide adaptive neighbor weighting, they lack several components crucial for handling the distinctive sparsity patterns and irregular geometry of LiDAR point clouds. Our preliminary testing and previous research [awedat2025] demonstrated that a single standard layer lacks the representational capacity required for high-fidelity LiDAR reconstruction. Consequently, our earlier work [awedat2025] relied on stacking multiple GAT layers to achieve competitive results. To address these limitations, we propose SuperiorGAT a specialized architecture with targeted enhancements for reconstructing missing points in sparse 3D scans. The overall framework comparison is illustrated in Figure 3(b), while the detailed computational pipeline is shown in Figure 4. By incorporating a gated residual connection, the model can preserve the original spatial coordinates (identity) while selectively integrating features learned through the attention mechanism. This approach allows SuperiorGAT to outperform multi-layer baselines using only a single refined layer, providing a more efficient solution for spatial data recovery.
@@
All experiments were conducted using the KITTI raw LiDAR dataset [Geiger2013IJRR], focusing on a residential driving sequence recorded with a Velodyne HDL-64E sensor operating at 10 Hz. The selected sequence contains 107 frames and represents a typical urban residential scenario. Each LiDAR scan provides three-dimensional point coordinates, from which approximately 50,000 points per frame were retained using stratified sampling to preserve the vertical beam structure.
To simulate realistic sensor degradation, structured beam dropout was applied by removing every fourth LiDAR beam, resulting in approximately 25% data loss per frame. This process introduces vertical sparsity that reflects common failure modes in cost-effective or degraded LiDAR systems. The dropout pattern was fixed and applied consistently across frames.
Each sparse point cloud was represented as a graph using a $k$-nearest neighbor construction, where $k$ denotes the neighborhood size. Node features consisted of the spatial coordinates $(x,y)$, a masked $z$-coordinate (set to zero for dropped points), and the corresponding beam index. Graphs were constructed independently for each frame.
To isolate the effectiveness of the architectural refinement, models were trained and evaluated independently per frame. Following the constraints established in our previous work [awedat2025], no temporal information, RGB data, or cross-frame learning was utilized in any experiment, ensuring the reconstruction relies purely on geometric graph dependencies within a single sweep..
All models were implemented using PyTorch and PyTorch Geometric. Performance was evaluated exclusively on the dropped points using root mean square error (RMSE) and Chamfer distance, along with per-frame inference time to assess computational efficiency.
@@
The performance of the SuperiorGAT model was first evaluated through a sensitivity analysis of the neighborhood parameter $k$ using the residential dataset. This parameter is critical as it defines the local graph connectivity and directly influences the receptive field of the attention mechanism. To identify the optimal balance between reconstruction accuracy and computational efficiency, we varied $k\in\{5,10,15,20\}$, with the results illustrated in Fig. 5.

### v3

@@ Abstract
LiDAR-based perception in autonomous systems is fundamentally constrained by fixed vertical beam resolution and is further degraded by structured beam dropout caused by occlusions or reduced-cost sensing hardware. This paper introduces SuperiorGAT, a graph attention–based framework for reconstructing missing elevation information in sparse LiDAR point clouds under structured beam loss. By modeling LiDAR scans as beam-aware graphs and augmenting standard graph attention networks with gated residual fusion and lightweight feed-forward refinement, the proposed approach improves vertical reconstruction accuracy without increasing network depth.
The effectiveness of SuperiorGAT is evaluated through extensive experiments on multiple KITTI environments, including Person, Road, Campus, and City, as well as cross-dataset validation on nuScenes with lower vertical resolution. Robustness is assessed under increasing structured beam dropout, demonstrating that SuperiorGAT consistently achieves lower reconstruction error and improved geometric consistency compared to interpolation-based methods, PointNet-based models, and deeper GAT baselines. Qualitative X–Z projection analyses further confirm the model’s ability to preserve structural continuity with minimal vertical distortion. Overall, the results indicate that targeted architectural refinement provides a computationally efficient solution for enhancing LiDAR vertical reconstruction without reliance on additional sensor modalities or hardware upgrades.
@@ Introduction
Autonomous vehicles rely critically on Light Detection and Ranging (LiDAR) sensors for 3D environmental perception, enabling precise object detection, localization, and mapping essential for safe navigation [29, 11]. Despite their fundamental importance, LiDAR systems face a critical vulnerability: performance severely degrades under structured data loss, particularly beam dropout caused by hardware malfunctions or environmental occlusions, which disrupts spatial continuity and compromises downstream tasks such as object detection and path planning [29, 3]. This challenge is particularly acute for cost-effective LiDAR systems that employ fewer scanning beams (typically 16-32 beams compared to 64+ in premium systems), inherently producing sparse point clouds that demand robust reconstruction methods to achieve the high-resolution performance required for autonomous operation [8].
@@
Current reconstruction approaches face a fundamental trade-off between accuracy and computational efficiency, limiting their practical deployment. Compressive sensing (CS) methods provide theoretical guarantees for sparse signal recovery but depend on computationally intensive iterative solvers that render them impractical for the real-time constraints of autonomous systems [6, 7]. Convolutional neural networks (CNNs), particularly voxel-based architectures, demonstrate effectiveness in processing point clouds but necessitate dense 3D projections, resulting in inefficient memory usage and computational overhead when applied to inherently sparse LiDAR data [24, 16].
@@
Recent advances in deep learning, including hierarchical point cloud models that operate directly on unordered point sets, have shown promise but remain inadequate for addressing the structured sparsity patterns characteristic of beam dropout scenarios [19]. While Graph Neural Networks (GNNs), especially Graph Attention Networks (GATs), have demonstrated significant potential in modeling complex geometric relationships within irregular data structures through sophisticated graph representations [21, 23, 22], standard graph-based approaches, including graph convolutional networks, often fail to effectively prioritize the critical local geometric features essential for high-fidelity reconstruction of sparse LiDAR data [17]. Our recent conference work demonstrated that multi-layer Graph Attention Networks can effectively reconstruct missing LiDAR elevation values using raw point-cloud geometry alone, without relying on RGB images or temporal information [1]. While increasing network depth improved reconstruction fidelity, deeper attention stacks introduced higher computational cost and stability limitations. In this work, we move beyond depth-based improvements and focus on architectural refinement to achieve superior performance with a lightweight design.
@@
To address these fundamental limitations, we propose SuperiorGAT, a novel graph attention-based framework specifically engineered to reconstruct sparse LiDAR point clouds with exceptional fidelity while maintaining the computational efficiency necessary for real-time deployment in autonomous systems [16, 22]. Our approach directly tackles the structured sparsity challenge through several key innovations. SuperiorGAT implements a realistic beam dropout simulation framework that accurately mimics hardware faults observed in real-world LiDAR deployments, providing a robust testing environment for autonomous vehicle applications [29, 3]. The system constructs optimized k-nearest neighbor graphs to effectively represent sparse point cloud structures while incorporating specialized beam index features that enhance both spatial and structural contextual understanding [22].
@@
The model architecture features a carefully designed lightweight framework incorporating multiple graph attention layers with multi-head attention mechanisms and residual connections, enabling highly efficient processing of sparse data without the computational burden of iterative CS optimization or the memory overhead of dense CNN projections [23]. By directly addressing structured sparsity patterns, SuperiorGAT provides a scalable and practical solution for autonomous driving applications, successfully bridging the critical gap between high-resolution reconstruction accuracy and real-time computational efficiency. Our primary contributions to the field include:
@@
Light Detection and Ranging (LiDAR) technology is a cornerstone of 3D perception for autonomous vehicles, enabling tasks like object detection and mapping through precise geometric data [11, 3]. Yet, sparse point clouds—caused by low-resolution sensors or beam dropout from environmental factors such as rain or hardware faults—complicate accurate reconstruction [zhang2021]. As illustrated in Figure 1, beam dropout predominantly disrupts elevation (z-coordinates), while x and y coordinates, derived from azimuthal scans, remain stable, underscoring the need for targeted z-reconstruction methods [10, 14]. This vertical sparsity stems from the limited number of laser channels in affordable LiDAR systems, a challenge evident in datasets like KITTI [11].
@@
Conventional reconstruction techniques include interpolation and compressive sensing (CS). Interpolation methods, such as linear or nearest-neighbor approaches, estimate missing points based on proximity but often smooth out critical geometric features, particularly in complex scenes [6, 7, 20]. Advanced techniques like spline interpolation reduce this issue but can introduce artifacts in regions with sharp elevation shifts [20]. CS offers a theoretical framework for recovering sparse signals but demands computationally intensive iterative processes, rendering it unsuitable for real-time applications where sub-100ms inference is critical [6, 7, 23]. These limitations highlight the demand for efficient solutions tailored to LiDAR’s structured sparsity.
@@
Deep learning has transformed point cloud processing, with Convolutional Neural Networks (CNNs) like VoxelNet converting sparse data into dense voxel grids for tasks such as 3D object detection [24]. However, this densification escalates computational costs and erodes fine details in sparse LiDAR data [24, 28]. VoxelNet’s cubic scaling with resolution contrasts with PointNet++, which uses hierarchical sampling to manage irregular data, yet both falter with structured dropout patterns due to static feature aggregation [19, 15]. PointNet and its extensions leverage permutation invariance but struggle to adapt to the specific sparsity of beam dropout [19, 10].
@@
Recent advances in LiDAR super-resolution target vertical resolution enhancement, focusing on z-coordinates for improved pedestrian detection [28, 10]. You and Kim (2023) introduced a 2D range image-based upsampling technique, interpolating x, y, z, and intensity to boost 3D detection accuracy, while Eskandar et al. (2023) proposed HALS, a height-aware model using polar coordinate regression to excel on KITTI data [28, 10]. These approaches affirm z-reconstruction’s importance, given the reliability of x and y coordinates in LiDAR scans [zhang2021], though their reliance on dense architectures limits real-time viability.
@@
Graph Neural Networks (GNNs) provide a versatile framework for irregular data, modeling points as nodes and relationships as edges [21]. Graph Convolutional Networks (GCNs) apply fixed-weight convolutions, effective for node classification but less adept at prioritizing sparse LiDAR geometries [17]. Enhanced GNNs incorporate edge attributes like distances, enhancing spatial modeling [25, 22]. Graph Attention Networks (GATs), pioneered by Veličković et al, introduce dynamic weighting via attention mechanisms, drawing from neural machine translation insights [23, 2]. Though successful in network analysis, GATs remain underexplored for LiDAR, where their adaptive aggregation suits sparse geometries [25]. Recent work in graph neural networks has explored architectural elements that enhance stability and expressive capacity beyond the standard GAT formulation. In particular, gated skip connections have been shown to improve gradient flow and help models balance raw node features with attention-refined representations [18, 4]. In our prior work [1], we applied a multi-layer GAT architecture to LiDAR beam reconstruction and showed that attention-based message passing is effective for recovering missing elevation data. The present work extends this direction by introducing gated residual fusion and feed-forward refinement, enabling improved reconstruction without increasing graph depth. Likewise, several graph transformer variants incorporate feed-forward sublayers similar to those used in sequence transformers to increase nonlinear capacity and improve local geometric modeling [9]. These ideas motivate our adoption of a lightweight gated residual pathway and a transformer-style feed-forward block, adapted specifically to the characteristics of sparse LiDAR reconstruction.
@@
SuperiorGAT advances GATs by optimizing for LiDAR z-reconstruction, integrating beam index features to reflect scanning patterns and employing a lightweight design with multi-head attention and residuals [23, 25]. Unlike previous methods, it tackles structured beam dropout (Figure 1) with a focus on z-coordinates, offering a balance of accuracy and efficiency for autonomous driving. This sets it apart from dense CNN approaches [24], PointNet variants [19], and super-resolution techniques [28, 10], establishing a novel real-time solution.
@@
Graph Neural Networks (GNNs) represent a powerful class of deep learning models designed to perform inference on data described by graphs. By leveraging the inherent relational structure of graph data, GNNs overcome a key limitation of traditional neural networks, which require input data to exist in an independent and identically distributed (i.i.d.) or grid-like structure (e.g., images) [21, 27].
The core operational principle of GNNs is message passing, a framework where nodes in a graph iteratively aggregate information from their local neighbors to build increasingly sophisticated representations of themselves and their context [13, 27]. This process can be broken down into three steps at each layer $l$:
@@
This mechanism allows each node to progressively incorporate information from multi-hop neighborhoods. For example, in a 2-layer GNN, a central node aggregates features from its direct neighbors in the first layer, and from its neighbors’ neighbors in the second layer. This ability to capture higher-order structural dependencies makes GNNs particularly effective for tasks where node states are influenced by spatial and relational context [17, velickovic2018gat, 26].
@@
This graph structure enables the application of Graph Neural Networks (GNNs) to model local relationships among points, leveraging the beam-aware connectivity to infer missing z-values. The representation provides a foundation for advanced graph-based techniques, setting the stage for exploring convolutional and attention-based approaches to enhance reconstruction accuracy. This prepares the framework for subsequent development of tailored GNN architectures.
@@
Graph Attention Networks (GATs) represent a significant advancement in graph representation learning by introducing an attention mechanism that enables dynamic, content-aware neighborhood aggregation [23]. Unlike Graph Convolutional Networks (GCNs) that employ fixed-weight aggregation based on graph structure, GATs compute adaptive attention coefficients that weight the contribution of each neighbor during feature propagation, allowing the model to focus on the most relevant contextual information for sparse LiDAR reconstruction.
@@
$\alpha_{ij}=\frac{\exp(e_{ij})}{\sum_{k\in\mathcal{N}(i)}\exp(e_{ik})}$ (6)
@@
where $\sigma$ is a nonlinear activation function. This formulation enables anisotropic aggregation, where each neighbor’s contribution is dynamically weighted based on feature compatibility, providing a more expressive scheme than isotropic approaches [23, 13]. The standard GAT architecture follows the sequential processing pipeline illustrated in Figure 3a.
@@
While standard GATs provide adaptive neighbor weighting, they lack several components crucial for handling the distinctive sparsity patterns and irregular geometry of LiDAR point clouds. To address these limitations, we propose SuperiorGAT—a specialized architecture with targeted enhancements for reconstructing missing points in sparse 3D scans. The overall framework comparison is illustrated in Figure 3, while the detailed computational pipeline is shown in Figure 4.
@@
All experiments were conducted using two public autonomous driving LiDAR datasets: the KITTI raw dataset [12] and the nuScenes dataset [5]. These datasets were selected to evaluate the proposed method under different sensor configurations and scene characteristics.
For the KITTI dataset, experiments focused on a residential driving sequence recorded with a Velodyne HDL-64E sensor operating at 10 Hz. The selected sequence contains 200 frames and represents a typical urban residential scenario. Each LiDAR scan provides three-dimensional point coordinates, from which approximately 50,000 points per frame were retained using stratified sampling to preserve the vertical beam structure.
In addition to the KITTI dataset, we evaluated the proposed framework on the nuScenes dataset [5] to assess robustness across LiDAR sensors with differing vertical resolutions. nuScenes is a large-scale autonomous driving dataset collected in urban environments using a Velodyne HDL-32E LiDAR sensor operating at 20 Hz, which provides 32 vertical scanning beams, half the vertical resolution of the KITTI sensor.
For nuScenes experiments, LiDAR frames were processed using the same pipeline applied to KITTI, with beam-aware graph construction and per-frame independent reconstruction. Due to the reduced number of vertical beams, structured beam dropout was applied proportionally by removing fixed subsets of beams to simulate degradation rates of 12.5%, 25%, and 37.5%. This setting reflects realistic failure or sparsification scenarios commonly encountered in lower-cost LiDAR systems.
All models were evaluated using identical graph parameters and architectural configurations to those used in the KITTI experiments, including the same neighborhood size $k$. Performance metrics were computed exclusively on the dropped points using RMSE in the vertical direction and geometric consistency measures. This cross-dataset evaluation allows direct comparison of reconstruction behavior under varying sensor resolutions while maintaining a consistent experimental protocol. Models are trained and evaluated independently on each dataset to ensure fair comparison under dataset-specific sensor configurations, without cross-dataset fine-tuning.
@@
Neighborhood Sensitivity Analysis
We first analyze the sensitivity of the proposed SuperiorGAT model to the neighborhood size parameter $k$, which determines the number of neighboring points used to construct the local graph for attention-based aggregation. This parameter directly controls the balance between local geometric detail and broader spatial context. The analysis is conducted on the KITTI Residential dataset, and the results are illustrated in Fig. 5.